OpenStudy (anonymous):

What are the approximate solutions of 2x² – x + 10 = 0? –2, 2.5 –1.97, 2.47 –2.5, 2 There is no solution.

OpenStudy (anonymous):

2x^2-x+10=0 2x^2-x+10-10=0-10 2x^2-x-10 2x^2-x+x=-10+x 2x^2=-10+x there is no solution

OpenStudy (anonymous):

thanks